Trigonometry
Solve each equation; clearly state the principle root, all solutions in [0,2pie],
and allreal solutions. If a calculator is necessary, round to four decimal places.
1) 9 cos x = 4
2) root 2 csc x+3=7
3) root 3 sin x tan (2x) - sin x=0

Answer by lwsshak3(11628)

Solve each equation; clearly state the principle root, all solutions in [0,2pie],
and allreal solutions. If a calculator is necessary, round to four decimal places.
1) 9 cos x = 4
cosx=4/9
x=1.1102, 5.1729
..
2) root 2 csc x+3=7
csc x=4/√2=1/sinx
sinx=√2/4
x=0.3614, 2.7802
..
3) root 3 sin x tan (2x) - sin x=0
sinx(√3tan2x-1)=0
sinx=0
x=0, π
or
√3tan2x-1=0
tan2x=1/√3
2x=π/6, 7π/6
x=π/12, 7π/12